# Semantic Edge Computing and Semantic Communications in 6G Networks: A Unifying Survey and Research Challenges

### v2

@@ 3.3.2 Bit-Rate Reduction
Source coding is an effective way to compress the information in original data with fewer bits. For example, Carra and Neglia [45] argued that a carefully designed linear quantization and run-length coding approach can achieve similar compression performance in comparison to bottleneck approaches. Therefore, various bit-rate reduction technique are leveraged in literature for semantic compression. Based on the source coding technique, literature can be categorized with lossless coding (e.g. run-length coding, entropy coding) and lossy coding (e.g., JPEG, quantization). For example, Chen et al. [83] benchmarked various existing lossless (e.g. GZIP [93]; ZLIB [94]; BZIP2 [95]; and LZMA [96]) and lossy compression (HEVC [97]) techniques on intermediate representations for different DNN architectures. Ko et al. [81] applied run-length and Huffman coding to achieve 3-10x compression rate and JPEG to achieve up to 50x compression rate without bottleneck. However, simply applying lossless coding can achieve limited compression performance. To achieve a significant compression rate, Cohen et al. [82] introduced a light-weight compression method for pre-trained object detectors based on clipping, binarization and entropy coding. Wang et al. [84] proposed a multi-stage quantization algorithm based on feature statistics and greedy search to minimize the information loss in bit-rate reduction. Alvar and Bajić [85] studied the optimal bit allocation of intermediate feature for multi-stream tasks by modeling it as a multi-objective optimization problem.
Note that bit-rate reduction can be applied on top of dimension reduction approaches such as bottleneck compression [73, 74, 6]. For example, Eshratifar et al. [4] used JPEG to further compress the latent features after dimension reduction for image classification tasks. As the JPEG compressor is non-differentiable, an identical mapping is used to replace the compressor during backpropagation in order to train the model. Lee et al. [46] leveraged binarization and PNG encoding to reduce the size of bottleneck output with only 1% performance loss for object detection tasks compared to YOLOv5 [98]. Li et al. [80] first leveraged channel-wise attention [99] to prune the auto-encoder and then applied entropy coding to reduce the bit-rate of the compressed semantic. Guo et al. [79] proposed to use pruning to reduce the dimension and a learnable non-linear quantization block [100] to compress the semantic with an adaptive bit width.
One challenge of bit-rate reduction lies in the extra complexity introduced by the source coding algorithm. To find the optimal bit-rate allocation, literature usually leverage complex optimization approaches which introduce additional latency during inference [85, 84, 45]. On the other hand, simple coding scheme such as JPEG and uniform bit quantization may result in information loss and non-optimal compression rate [86].
@@ 3.3.3 Rate-Distortion Optimization
Inspired by the rate-distortion optimization, Ballé et al. [47] first proposed a neural image compression method based on a variational auto-encoder [102] and a factorized prior. A general framework of learning-driven entropy coding is depicted as Figure 7. A variational auto-encoder is used to compress and quantize the image to latent representation while a prior model is used to estimate the entropy of the compressed information. According to Shannon’s Coding Theorem [101], the lossless coding rate of a message is lower bounded by its entropy. Therefore, the objective of the image compression is to optimize the following rate-distortion trade-off:
@@
While entropy-based optimization approach is originally proposed for image compression, it is adopted to semantic compression rapidly. For example, [87] applied the entropy-based compression before the pre-trained DNN to directly compress the raw input. As the pre-trained DNN is not optimized for reconstructed input, it requires a multi-stage training. The first stage is to train the compressor with Eqn. 3 and the second stage is to fine-tune the DNN with task-oriented metrics (e.g. cross-entropy loss). In addition to add compression before pre-trained DNNs, other work injects entropy-based compression in the middle of DNNs. For example, in [88], an auto-encoder with factorized prior [47] is trained to reconstruct the latent representation of pre-trained DNNs. Datta et al. [89] proposed a similar approach but with a different entropy estimation module [104]. Ahuja et al. [90] injected the entropy compressor to compress the semantic and applied knowledge distillation to fine-tune the performance of edge DNN. Hossain et al. [91] does not optimize to reconstruct the latent representation but replace the distortion $D$ in Eqn. 3 with cross-entropy loss. To reduce the computation overhead introduced by entropy-based compression, Matsubara et al. [16] directly applied the rate-distortion optimization to exisiting light-weight bottleneck [6] and the multi-stage training approach in [6] is also applied to preserve the task-oriented performance. Matsubara et al. [86] systematically benchmarked the compression performance of [16] and other dimension or bit-rate reduction approaches. Experiments demonstrated that the learnable entropy coding can achieve the best compression-computation-performance trade-off in comparison to dimension reduction and naive bit-rate reduction.
@@
However, as the entropy estimation module is learned over a specific dataset, a distribution shift in data source may drastically affect the source coding rate. Zhang et al. [105] demonstrated that the compression performance of learned entropy model is vulnerable to interference such as short noise and adversarial attack. Therefore, the major challenge of entropy-based rate-distortion optimization is to train a compression module that is resilient to intentional and unintentional noise.
@@ 4.1 Taxonomy of Semantic Communication
The taxonomy of SemCom is shown in Figure 8. We classify recent advancements in SemCom into two categories: train-time static coding and test-time adaptive coding. Static coding focuses on generating semantic representations that are resilient to stochastic wireless channels, and can be further divided into semantic-level coding and task-level coding. On the other hand, adaptive coding aims to generate semantics dynamically in response to varying channel conditions, which can be further categorized into channel-adaptive coding and importance-aware coding. Table 3 summarizes all the literature reviewed in this paper.
@@ 4.1.1 Train-Time Static Systems
Early work [106, 23] in this domain assumed a simple Additive White Gaussian Noise (AWGN) model for the wireless channel which is implemented using a non-trainable noise-injection layer and added between the encoder and decoder DNNs during training. Such a model fails to express the real-world channel characteristics which in turn limits the performance of such SemCom systems after deployment. Other work [24, 107, 108] proposed different semantic similarity metrics or more expressive architectures for the DNN to help improve the trained system generalization ability to real-world channels. To support multi-path fading effects, Felix et al. [109], O’Shea et al. [110] employ Raleigh and Rician models for channel impairments. However, due to the mismatch between the stochastic channel model used for training and the actual channel transfer function during test time, real-world implementations suffer from a noticeable performance drop compared to the results obtained during simulation.
@@ 4.1.2 Test-Time Adaptive Systems
While in these methods the architecture of the encoder and decoder DNNs are fixed and only their parameters are updated, recent work [48] use dynamic pruning to change the number of transmitted symbols based on the current wireless link quality. Moreover, other work [116] extract the importance score of different semantics and prioritize the transmission of the most important ones in severe channel conditions.
@@ 4.2.1 Semantic-Level Coding
O’shea and Hoydis [106] first proposed using an auto-encoder with a noise-injection layer to implement an end-to-end communication system assuming an AWGN channel model. More specifically, the encoder and decoder part of the auto-encoder are deployed at the transmitter and receiver respectively, while the non-trainable noise injection layer is employed only during training for the DNN to learn the channel distortion patterns. O’Shea et al. [117] extended this idea to account for more realistic channel models such as Rayleigh multi-path fading. Also, Erpek et al. [21] improved the study by developing a Physical Layer (PHY) scheme for Multi-Input-Multi-Output (MIMO) systems in the presence of interference. All of these methods obtained superior performance compared to classical modular communication system designs.
Text transmission. The aforementioned DL-based communication systems laid the foundation for the development of SemCom. As such, Long Short-term Memorys (LSTMs) are used in [22] to perform JSCC for text transmission. The proposed method is capable of preserving semantic information by mapping the semantically-similar sentences onto closer points in the embedding space and using the Word Error Rate (WER) instead of Bit Error Rate (BER) to train the DNN. Having a similar design to the second level of Figure 2, Xie et al. [24] (DeepSC) used transformers [118] both as the semantic encoder and the decoder and added dense layers on top of them for channel coding which facilitates subsequent transmission. They developed a loss function $L=CE(S,\hat{S})-\lambda I(X;Y)$ to train the entire system, where the first term minimizes the cross-entropy (semantic difference) between $S$ and $\hat{S}$ and the second term maximizes robustness to channel distortions. As illustrated in Figure 9, the training framework of DeepSC has two stages: In phase 1, the mutual information estimation model is trained, and in phase 2, the whole system is trained considering both the cross entropy and the mutual information loss.
@@
A lightweight version of DeepSC is developed in [108], where the authors pruned the DNN and reduced the resolution of its parameters to obtain an IoT-friendly model. Sana and Strinati [119] adopted a novel loss function $L=I(X;S)-(1+\alpha)I(X;Y)+\beta KL(Y,\hat{S})$ by taking into account the compactness of the extracted semantics. In this loss function, the first two terms are used to ensure compactness and robustness respectively, while the last term minimizes the semantic distortion incurred during transmission. Moreover, the authors proposed a semantic-adaptive mechanism which dynamically adapts the number of symbols per word to achieve a better trade-off between performance and communication efficiency.
@@
Image transmission. JSCC schemes have been employed for image transmission which directly map raw pixel values to complex-valued channel symbols. In these schemes, the DL-based semantic encoder first extracts low-dimensional semantic representations which are then distorted according to a channel model. The semantic decoder recovers the images based on the received symbols, where the end-to-end system is trained to maximize the semantic similarity metrics between the original and recovered images such as Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Method (SSIM). Bourtsoulatze et al. [23] first proposed a technique to transmit images called DeepJSCC, where a Convolutional Neural Network (CNN)-based autoencoder is used for JSCC. Kurka and Gündüz [107] introduced the variable-length coding feature by utilizing channel feedback (i.e., DeepJSCC-f), where the transmission of each image is done sequentially in $L$ stages. Specifically, each stage improves the quality of the recovered image by sending additional semantics corresponding to the residual error from the previous stage. Kurka and Gündüz [120] redesigned a DeepJSCC framework for adaptive-bandwidth transmission, where images are transmitted progressively in layers over time. As such, they proposed a successive refinement scheme for JSCC where the receiver can aggregate the received layers sequentially to increase the quality of the reconstructed semantics. Zhang et al. [121] proposed a multi-level semantic feature extractor for image transmission, called MLSC-image, which encodes both the high-level information such as text and segmentation semantics and the low-level information such as local spatial details. These extracted features are later combined and encoded into channel symbols.
@@
While the aforementioned work study the optimization of differentiable objectives, other methods focus on utilizing DRL to address user-defined non-differentiable semantic metrics which commonly arise in real-world wireless systems. Lu et al. [25] first proposed a DRL-based SemCom system, called SemanticRL-JSCC, where a self-critic policy is learned to maximize the semantic similarity measures such as Bilingual Evaluation Understudy (BLEU) [122] and Consensus-based Image Description Evaluation (CIDEr) [123]. Furthermore, to address non-differentiable channels, SemanticRL-JSCC learns separate policies for the encoder and decoder which shows strong robustness under fluctuating Signal-to-Noise Ratios (SNRs). Lu et al. [124] also used DRL to develop a joint semantics-noise coding (JSNC) method capable of adapting to time-varying channels while ensuring the preservation of semantic information. As such, a confidence-based distillation mechanism is employed in [124] both at the encoder and decoder which ensures the confidence reaches a threshold before the encoder/decoder could further process the semantics. Huang et al. [125] proposed a RL-based adaptive semantic coding (RL-ASC) technique which allocates different quantization levels for different semantic concepts based on their importance. Specifically, the crucial semantics are encoded with a higher precision to minimize a triple-term (rate, semantic, and perceptual) loss, while the task-irrelevant concepts can be discarded. Apart from the RL-based bit allocation module, [125] used a convolutional semantic encoder, and a GAN-based decoder equipped with an attention module which can fuse both local and global features.

### v3

@@ 3.3.2 Bit-Rate Reduction
Source coding is an effective way to compress the information in original data with fewer bits. For example, Carra and Neglia [68] argued that a carefully designed linear quantization and run-length coding approach can achieve similar compression performance in comparison to bottleneck approaches. Therefore, various bit-rate reduction technique are leveraged in literature for semantic compression. Based on the source coding technique, literature can be categorized with lossless coding (e.g. run-length coding, entropy coding) and lossy coding (e.g., JPEG, quantization). For example, Chen et al. [115] benchmarked various existing lossless (e.g. GZIP [127]; ZLIB [128]; BZIP2 [129]; and LZMA [130]) and lossy compression (HEVC [131]) techniques on intermediate representations for different DNN architectures. Ko et al. [113] applied run-length and Huffman coding to achieve 3-10x compression rate and JPEG to achieve up to 50x compression rate without bottleneck. However, simply applying lossless coding can achieve limited compression performance. To achieve a significant compression rate, Cohen et al. [114] introduced a light-weight compression method for pre-trained object detectors based on clipping, binarization and entropy coding. Wang et al. [116] proposed a multi-stage quantization algorithm based on feature statistics and greedy search to minimize the information loss in bit-rate reduction. Alvar and Bajić [117] studied the optimal bit allocation of intermediate feature for multi-stream tasks by modeling it as a multi-objective optimization problem. Xiao et al. [118] proposed a block-wise variable bit-rate quantization algorithm based on the attention map weight in transformers.
Note that bit-rate reduction can be applied on top of dimension reduction approaches such as bottleneck compression [103, 104, 6]. For example, Eshratifar et al. [4] used JPEG to further compress the latent features after dimension reduction for image classification tasks. As the JPEG compressor is non-differentiable, an identical mapping is used to replace the compressor during backpropagation in order to train the model. Lee et al. [69] leveraged binarization and PNG encoding to reduce the size of bottleneck output with only 1% performance loss for object detection tasks compared to YOLOv5 [132]. Li et al. [112] first leveraged channel-wise attention [133] to prune the auto-encoder and then applied entropy coding to reduce the bit-rate of the compressed semantic. Guo et al. [110] proposed to use pruning to reduce the dimension and a learnable non-linear quantization block [134] to compress the semantic with an adaptive bit width. Li et al. [119] used auto-encoder to compress the dimension and applied a variable depth quantization based on the attention map of CNN channels. Entropy coding is further leveraged to compress the datasize.
@@ 3.3.3 Rate-Distortion Optimization
Inspired by the rate-distortion optimization, Ballé et al. [70] first proposed a neural image compression method based on a Variational Auto-Encoder (VAE) [136] and a factorized prior. A general framework of learning-driven entropy coding is depicted as Figure 7. A VAE is used to compress and quantize the image to latent representation while a prior model is used to estimate the entropy of the compressed information. According to Shannon’s Coding Theorem [135], the lossless coding rate of a message is lower bounded by its entropy. Therefore, the objective of the image compression is to optimize the following rate-distortion trade-off:
@@
While entropy-based optimization approach is originally proposed for image compression, it is adopted to semantic compression rapidly. For example, [121] applied the entropy-based compression before the pre-trained DNN to directly compress the raw input. As the pre-trained DNN is not optimized for reconstructed input, it requires a multi-stage training. The first stage is to train the compressor with Eqn. 3 and the second stage is to fine-tune the DNN with task-oriented metrics (e.g. cross-entropy loss). In addition to add compression before pre-trained DNNs, other work injects entropy-based compression in the middle of DNNs. For example, in [123], an auto-encoder with factorized prior [70] is trained to reconstruct the latent representation of pre-trained DNNs. Datta et al. [124] proposed a similar approach but with a different entropy estimation module [138]. Ahuja et al. [125] injected the entropy compressor to compress the semantic and applied knowledge distillation to fine-tune the performance of edge DNN. Hossain et al. [61] does not optimize to reconstruct the latent representation but replace the distortion $D$ in Eqn. 3 with cross-entropy loss. To reduce the computation overhead introduced by entropy-based compression, Matsubara et al. [58] directly applied the rate-distortion optimization to exisiting light-weight bottleneck [6] and the multi-stage training approach in [6] is also applied to preserve the task-oriented performance. Matsubara et al. [120] systematically benchmarked the compression performance of [58] and other dimension or bit-rate reduction approaches. Experiments demonstrated that the learnable entropy coding can achieve the best compression-computation-performance trade-off in comparison to dimension reduction and naive bit-rate reduction. Harell et al. [122] systematically investigated the theory and designing strategies including the splitting point choice, training strategy and architecture design of learning-driven rate-distortion optimization in split computing. Yuan et al. [126] created layers with different dimensions to achieve an adaptive compression performance and proposed a multi-round refinement method to optimize the rate-distortion.
3.3.4 Research Challenges
Bottleneck Optimization. Injecting bottleneck in DNNs to reduce data dimension can jeopardize the task-oriented performance. To preserve the accuracy, research usually involves advanced training methods to optimize the bottleneck. For example, Yao et al. [103] applied multi-stage training to auto-encoder. In the first stage, the auto-encoder is trained to reconstruct the semantic features and in the second stage, knowledge distillation is applied to fine-tune the auto-encoder. Hao et al. [104] add a performance regularization term to reconstruction loss when training the auto-encoder. After training the auto-encoder, the partitioned DNN on the edge device is fine-tuned to further recover the performance. Matsubara et al. [6] first applied knowledge distillation to train the encoder and then fine-tune the decoder to further mitigate the performance drop.
Source Coding Complexity. To achieve better data compression performance, literature usually leverage complex optimization approaches to find the optimal bit-rate allocation for different semantic features [117, 116, 68]. However, optimization-based source coding can introduce additional latency during inference. On the other hand, simple coding scheme such as JPEG and uniform bit quantization may result in information loss and less optimal compression rate compared to others [120].
@@ 3.3.4 Research Challenges
Resilience of Entropy Models. Learning-driven rate-distortion methods can jointly optimize the bottleneck and source coding [125, 122], resulting in a better compression performance compared to other methods separately optimize bottleneck and source coding [6, 4]. However, as the entropy estimation module is learned over a specific dataset, a distribution shift in data source may drastically affect the source coding rate. Zhang et al. [139] demonstrated that the compression performance of learned entropy model is vulnerable to interference such as short noise and adversarial attack. Therefore, the major challenge of entropy-based rate-distortion optimization is to train a compression module that is resilient to intentional and unintentional noise.
@@ 4.1 Taxonomy of Semantic Communication
The taxonomy of SemCom is shown in Figure 8. We classify recent advancements in SemCom into two categories: train-time static coding and test-time adaptive coding. Static SemCom systems focus on generating semantic representations that are resilient to stochastic wireless channels, and can be further divided into semantic-level coding and task-level coding. On the other hand, adaptive SemCom systems aim to generate semantics dynamically in response to varying channel conditions, which can be further categorized into channel-adaptive coding and importance-aware coding. Table 4 summarizes the recent literature reviewed in this paper.
@@ 4.1.1 Train-Time Static Systems
Early work [140, 38] in this domain assumed a simple Additive White Gaussian Noise (AWGN) model for the wireless channel which is implemented using a non-trainable noise-injection layer and added between the encoder and decoder DNNs during training. Such a model fails to express the real-world channel characteristics which in turn limits the performance of such SemCom systems after deployment. Other work [39, 141, 142, 143] proposed different semantic similarity metrics or more expressive architectures for the DNN to help improve the trained system generalization ability to real-world channels. To support multi-path fading effects, Felix et al. [144], O’Shea et al. [145] employ Raleigh and Rician models for channel impairments. However, due to the mismatch between the stochastic channel model used for training and the actual channel transfer function during test time, real-world implementations suffer from a noticeable performance drop compared to the results obtained during simulation.
@@ 4.1.2 Test-Time Adaptive Systems
While in these methods the architecture of the encoder and decoder DNNs are fixed and only their parameters are updated, recent work [71] use dynamic pruning to change the number of transmitted symbols based on the ongoing wireless link quality. Moreover, other work [151] focuses on making the SemCom systems adaptive to the data rather than the channel. For example, the authors in [151] extract the importance score of different semantics and prioritize the transmission of the most important ones in severe channel conditions.
@@ 4.2.1 Semantic-Level Coding
O’shea and Hoydis [140] first proposed using an auto-encoder with a noise-injection layer to implement an end-to-end communication system assuming an AWGN channel model. More specifically, the encoder and decoder part of the auto-encoder are deployed at the transmitter and receiver, respectively, while the non-trainable noise injection layer is employed only during training for the DNN to learn the channel distortion patterns. O’Shea et al. [152] extended this idea to account for more realistic channel models such as Rayleigh multi-path fading. Also, Erpek et al. [36] improved the study by developing a Physical Layer (PHY) scheme for Multi-Input-Multi-Output (MIMO) systems in the presence of interference. All of these methods obtained superior performance compared to classical modular communication system designs.
@@
Text transmission. The aforementioned DL-based communication systems laid the foundation for the development of SemCom. As such, Long Short-term Memorys (LSTMs) are used in [37] to perform JSCC for text transmission. The proposed method is capable of preserving semantic information by mapping the semantically similar sentences onto closer points in the embedding space and using the Word Error Rate (WER) instead of Bit Error Rate (BER) to train the DNN. Having a similar design to the second level of Figure 2, Xie et al. [39] (DeepSC) used transformers [153] both as the semantic encoder and decoder and added dense layers on top of them for channel coding which facilitates subsequent transmission. They developed a loss function $L=CE(S,\hat{S})-\lambda I(X;Y)$ to train the entire system, where the first term minimizes the cross-entropy (semantic difference) between $S$ and $\hat{S}$ and the second term maximizes robustness to channel distortions. As illustrated in Figure 9, the training framework of DeepSC has two stages: In phase 1, the mutual information estimation model is trained, and in phase 2, the whole system is trained considering both the cross entropy and the mutual information loss.
@@
A lightweight version of DeepSC is developed in [142], where the authors prune the DNN and reduce the resolution of its parameters to obtain an Internet of Things (IoT)-friendly model. Sana and Strinati [154] adopted a novel loss function $L=I(X;S)-(1+\alpha)I(X;Y)+\beta KL(Y,\hat{S})$ by taking into account the compactness of the extracted semantics. In this loss function, the first two terms are used to ensure compactness and robustness respectively, while the last term minimizes the semantic distortion incurred during transmission. Moreover, the authors proposed a semantic-adaptive mechanism which dynamically adapts the number of symbols per word to achieve a better trade-off between performance and communication efficiency.
@@
Image transmission. JSCC schemes have been employed for image transmission which directly map raw pixel values to complex-valued channel symbols. In these schemes, the DL-based semantic encoder first extracts low-dimensional semantic representations which are then distorted according to a channel model. The semantic decoder recovers the images based on the received symbols, where the end-to-end system is trained to maximize the semantic similarity metrics between the original and recovered images such as Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Method (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) [155]. Bourtsoulatze et al. [38] first proposed a technique to transmit images called DeepJSCC, where a Convolutional Neural Network (CNN)-based autoencoder is used as the JSCC model. Kurka and Gündüz [141] introduced the variable-length coding feature by utilizing channel feedback (i.e., DeepJSCC-f), where the transmission of each image is done sequentially in $L$ stages. Specifically, each stage improves the quality of the recovered image by sending additional semantics corresponding to the residual error from the previous stage. Kurka and Gündüz [156] redesigned a DeepJSCC framework for bandwidth-adaptive transmission, where images are transmitted progressively in layers over time. As such, they proposed a successive refinement scheme for JSCC where the receiver can aggregate the received layers sequentially to increase the quality of the reconstructed semantics. Zhang et al. [157] proposed a multi-level semantic feature extractor for image transmission, called MLSC-image, which encodes both the high-level information such as text and segmentation semantics and the low-level information such as local spatial details. These extracted features are later combined and encoded into channel symbols. To make JSCC systems compatible with digital modulation schemes, Bo et al. [158] developed a joint coding-modulation framework, called JCM, based on VAEs that outputs discrete constellation symbols as the encoded representation. The presented framework is trained in an end-to-end fashion to match the obtained modulation strategy with the operating channel condition. The authors in [159] improved the robustness of digital SemCom systems by developing a demodulation method that can assess the uncertainty of the demodulation output. Moreover, they considered a binary symmetric erasure model for the channel and proposed adaptively determining the modulation order based on channel conditions.
Generative AI has been leveraged in [160] to increase the perceptual quality of the reconstructed images in SemCom systems. More specifically, the source and channel coding mechanism at the transmitter end is implemented by a Swin Transformer, while a diffusion model is used at the receiver to decode high-quality images. Also, Wu et al. [161] proposed using a diffusion model to gradually remove the channel noise in SemCom systems. To achieve this, the diffusion model is applied after the channel equalization module and learns the distribution of the channel input signal. Tang et al. [162] utilize contrastive learning to develop a semantic contrastive coding scheme. In the proposed scheme, input images and their reconstructed outputs are projected onto a latent space, where the projections of similar samples are pulled together, and the projections of irrelevant inputs are pushed apart following contrastive learning ideas. Moreover, when a DNN for the downstream task is not available, a duplicate of the semantic encoder is deployed at the receiver to facilitate the training process.
@@
While the aforementioned studies have differentiable objectives to optimize, other methods focus on utilizing DRL to address user-defined non-differentiable semantic metrics which commonly arise in real-world wireless systems. Lu et al. [40] first proposed a DRL-based SemCom system, called SemanticRL-JSCC, where a self-critic policy is learned to maximize the semantic similarity measures such as Bilingual Evaluation Understudy (BLEU) [163] and Consensus-based Image Description Evaluation (CIDEr) [164]. Furthermore, to address non-differentiable channels, SemanticRL-JSCC learns separate policies for the encoder and decoder which shows strong robustness under fluctuating Signal-to-Noise Ratios (SNRs). Lu et al. [165] also used DRL to develop a joint semantics-noise coding (JSNC) method capable of adapting to time-varying channels while ensuring the preservation of semantic information. As such, a confidence-based distillation mechanism is employed in [165] both at the encoder and decoder which ensures the confidence reaches a threshold before the encoder/decoder could further process the semantics. Huang et al. [166] proposed a RL-based adaptive semantic coding (RL-ASC) technique which allocates different quantization levels for different semantic concepts based on their importance. Specifically, the crucial semantics are encoded with a higher precision to minimize a triple-term (rate, semantic, and perceptual) loss, while the task-irrelevant concepts can be discarded. Apart from the RL-based bit allocation module, [166] used a convolutional semantic encoder, and a GAN-based decoder equipped with an attention module which can fuse both local and global features.
Some existing work utilizes multi-modal features to improve the transmission quality of SemCom systems. For example, Zhao et al. [167] propose a SemCom system (LaMoSC) which integrates both the visual and textual features extracted from input images using an incorporated pre-trained Large Language Model (LLM). The authors in [168] propose using a multi-modal language model to align the semantics in image, audio, video, and text data, and then utilize a personalized LLM as the knowledge base both at the transmitter to extract the semantics and at the receiver to recover the messages. Moreover, a conditional Generative Adversarial Network (cGAN) is employed for channel estimation based on the transmitted and received pilot signals.